### 16 Feb 2024 v1

# Fusing Neural and Physical: Augment Protein Conformation Sampling with Tractable Simulations

Abstract
The protein dynamics are common and important for their biological functions and properties, the study of which usually involves time-consuming molecular dynamics (MD) simulations in silico. Recently, generative models has been leveraged as a surrogate sampler to obtain conformation ensembles with orders of magnitude faster and without requiring any simulation data (a ”zero-shot” inference). However, being agnostic of the underlying energy landscape, the accuracy of such generative model may still be limited. In this work, we explore the few-shot setting of such pre-trained generative sampler which incorporates MD simulations in a tractable manner. Specifically, given a target protein of interest, we first acquire some seeding conformations from the pre-trained sampler followed by a number of physical simulations in parallel starting from these seeding samples. Then we fine-tuned the generative model using the simulation trajectories above to become a target-specific sampler. Experimental results demonstrated the superior performance of such few-shot conformation sampler at a tractable computational cost.
1 Introduction
Elucidating protein dynamics stands as a significant yet challenging problem during the study of protein functionality and regulation. Such macromolecule can go through transitions between multiple conformational states for different length or time scales. For example, the SARS-Cov-2 spike protein was found to have transitions between its open and closed states (Gur et al., 2020) for its function. Experimental instruments including crystallographic B-factors and NMR spectroscopy can be leveraged to probe such dynamics but only available to a limited spatial and temporal scale.
Traditionally, computational simulation methods especially molecular dynamics (MD) are used to simulate the dynamic behavior of these biological molecules. MD operates by simply evolving the Newtonian equation on the whole system of particles where the accelerations, or the gradients of velocity, are determined by a pre-specified force field (energy, or unnormalized density). To study the protein dynamics, the time scale of MD simulations can reach micro- to milli-seconds in order to completely capture the behaviors such as the transition from unfolding to folding (Lindorff-Larsen et al., 2011). However, simulating protein systems in such time scales is very computationally intensive, which can take several hundreds of GPU days , depending on the size of the system.
With the emergence of accurate structure prediction models (Jumper et al., 2021; Baek et al., 2021; Lin et al., 2023), deep generative models are developed to serve as efficient surrogate for MD simulations, such as Str2Str (Lu et al., 2024). Str2Str learns to explore the conformation space by training on protein structures from the Protein Data Bank (PDB) in an amortized way, and is ready to perform zero-shot conformation sampling for any unseen test proteins. The ability of direct sampling makes such sampler become orders of magnitude more efficient than traditional MD simulations. The inference of such diffusion sampler is based on successful transfer learning of the naturally occurred protein geometries. However, training solely on PDB data can make it fail to take care of the underlying energy landscape, which is, on the contrary, the fundamental consideration in MD simulation to generate Boltzmann-distributed ensembles.
In this work, we aim to improve the diffusion sampler with tractable MD simulations via fine-tuning, forming a ”few-shot” [^1] Here we define ”few-shot” in the sense that a limited quota of physical simulations for the test protein can be acquired, in contrast to the zero-shot setting where no simulation sample is present. setting of Str2Str (Lu et al., 2024). To be specific, given a target protein during inference, zero-shot samples are firstly generated from the pre-trained conformation sampler, followed by a number of ”roll-out”, i.e. short MD simulations respectively initialized by each sample, which is highly parallelizable and thus efficient. Then the production trajectories are leveraged to fine-tune the sampler ad hoc for the target protein and reinforce the sampling accuracy. Experimental results demonstrate that such process significantly improve the quality of conformation sampling, at the cost of tractable computations compared to conventional simulations from scratch.
Figure 1: Illustrative diagram of fine-tuning the pre-trained diffusion sampler Str2Str (Lu et al., 2024). Firstly, initial conformation samples of the target protein are generated from a pre-trained network $\mathcal{S}_{\theta}$ parameterized by $\theta$, followed by parallel MD simulations respectively for each sample. The production trajectories are leveraged to make a target-specific sampler $\mathcal{S}_{\theta^{*}}$ via fine-tuning.
2 Related work
Protein conformation sampling
Prior to this work, several methods have been proposed to sample protein conformations. Depending on the requirement of simulation data for training, these methods can be divided into two categories: (1) Require simulation data. Boltzmann generator (Noé et al., 2019) uses normalizing flow to approach the underlying Boltzmann distribution by learning from simulation data, which is one of the pioneer works among this research direction. Arts et al. (2023) instead adopts the denoising diffusion model to learn such distribution over the coarse-grained protein conformations. idpGAN (Janson et al., 2023) involves a conditional generative adversarial network trained on intrinsic disordered proteins. DiG (Zheng et al., 2023) trains a conditional diffusion model with both PDB and in-house simulation data and can sample conformations conditioned on input sequence. (2) Require no simulation data. AlphaFold2-RAVE (Vani et al., 2023) injects stochasticity to the structure output by making modifications to the input channel of AlphaFold2 (Jumper et al., 2021). EigenFold (Jing et al., 2023) revisits the structure prediction (folding) problem in a generative view and is able to sample from the distribution over conformations given the input sequence. Str2Str (Lu et al., 2024) is an equivariant diffusion sampler trained solely on PDB data and formulates the conformation sampling in a structure-to-structure manner.
De novo design of protein structure
Despite different purpose, protein structure design methods can also be relevant because of similar techniques used for modeling. Anand & Achim (2022) and Shi et al. (2022) leverage the backbone architecture of AlphaFold2 but feed forward from the input conditions such as contact map. Afterwards, generative models are applied to learn from PDB and de novo design novel backbone structures: ProtDiff (Trippe et al., 2022), FoldingDiff (Wu et al., 2022), RFDiffusion (Watson et al., 2023), Chroma (Ingraham et al., 2023), FrameDiff (Yim et al., 2023) and FoldFlow (Bose et al., 2023) to name a few. These sequence-agnostic generative models can design non-native backbone structures, based on which de novo sequences can further be designed by an inverse folding model, such as ProteinMPNN (Dauparas et al., 2022).
3 Methodology
3.1 Preliminaries
Protein structure.
Protein structure consists of a sequence of amino acids (or residues) which respectively defines a set of atoms existing in three-dimension (3D) space. Formally, a protein with $L$ residues, while the residue $i$ ($1\leq i\leq L$) is composed of $n_{i}$ atoms, can be represented by the Euclidean coordinates of atoms ${\mathbf{x}}\in\mathbb{R}^{N\times 3}$, where $N=\sum_{i=1}^{L}n_{i}$ is the totally number of atoms.
Score-based generative models.
Score-based generative models (SGMs) (Song et al., 2020) aim to model the data distribution via a diffusion process defined by the Itô stochastic differential equation (SDE):
${\textnormal{d}}{\mathbf{x}}={\mathbf{f}}({\mathbf{x}},t){{\textnormal{d}}}t+g(t){{\textnormal{d}}}{{\mathbf{w}}},$ (1)
where continuous time $t\in[0,T]$, ${\mathbf{f}}({\mathbf{x}},t)\in\mathbb{R}^{n}$ and $g(t)\in\mathbb{R}$. The ${\mathbf{w}}\equiv{\mathbf{w}}(t)\in\mathbb{R}^{n}$ represents the standard Wiener process. Then, the corresponding backward SDE can be expressed as follow  (Anderson, 1982):
${\textnormal{d}}{\mathbf{x}}=[{\mathbf{f}}({\mathbf{x}},t)-g^{2}(t)\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})]{\textnormal{d}}t+g(t){{\textnormal{d}}}{\bar{\mathbf{w}}},$ (2)
where ${\textnormal{d}}{t}$ is negative timestep and $\bar{{\mathbf{w}}}$ is the standard Wiener process with time $t$ flowing from $T\to 0$.
Zero-shot conformation sampler.
Str2Str (Lu et al., 2024) leveraged SGMs to sample protein backbone conformation via a perturbation followed by an annealing process. Str2Str was trained solely with the protein structures in PDB and requires no simulation data during both training and inference. Str2Str demonstrated the effectiveness of SGMs for conformation sampling in a transfer learning manner, which shows promise when long MD simulation data is not readily available.
Molecular dynamics simulation.
Molecular dynamics (MD) simulations amount to evolve the whole particle system through time $T^{\textrm{sim}}>0$  [^2] To avoid ambiguity with the time boundary $T$ in SGMs, the simulation time is denoted by $T^{\textrm{sim}}$ using superscript sim. by the physical dynamics ${d{\mathbf{x}}}={\mathbf{v}}({\mathbf{x}},\tau){d\tau}$, where the velocity field ${\mathbf{v}}\in\mathbb{R}^{N\times 3}$ can be determined by a pre-defined force field $U({\mathbf{x}})$ (aka energy or unnormalized density) via the Newtonian equation of motion: $d{\mathbf{v}}({\mathbf{x}},\tau)=\bar{{\mathbf{m}}}\nabla_{\mathbf{x}}U({\mathbf{x}})~{}d\tau$, where $\bar{{\mathbf{m}}}$ is the inverse mass of atoms in the system. At each time step, the position and velocity may be updated by a symplectic integrator such as Verlet (Verlet, 1967), yielding a simulation trajectory.
3.2 Neighborhood exploitation with short MD
The dilemma in direct sampling and simulation.
One disadvantage of Str2Str is that it has no prior knowledge about what do good conformations of the test protein look like during inference. On the other hand, MD simulations simply query a universal oracle (force field) at each time step, being generalizable and training-free, but have a hard time sufficiently exploring the conformation space within tractable computation for simulation. Previous methods, such as Arts et al. (2023), proposed to train using the simulation trajectory of a target protein and sample like a generative model. Although achieving good performance, such method however requires expensive simulation data beforehand and cannot quickly generalize to other proteins beyond the training target, which in practice may have restricted potential utility. Here arises an interesting question: how to organically combine the advantages of both generative models and physical simulations?
Figure 2: Illustration of the conformation sampling of Str2Str (Lu et al., 2024). In (a), sampling is performed in two steps to obtain i.i.d. samples of the target protein, where the energy landscape (or information of the force field) is unknown and colored gray. In (b), a hypothetical zoom-in neighborhood of a sample is shown. Due to the generally complicated landscape of force field, the samples directly generated by Str2Str are probably not energy-optimal. Short MD simulation can be run to obtain locally equilibrium samples and can be used to fine-tune the pre-trained sampler.
Exploration-exploitation balance.
To answer this question, we propose to sample protein ensembles via a two-stage sampling, or exploration-exploitation of the conformation space. Given a target protein of interest, we firstly leverage the Str2Str to sample a set of plausible conformations $X\equiv\{{\mathbf{x}}^{(i)}\}_{i=1}^{m}$, which can be seemingly good but not energy favorable. On top of that, a short MD simulation can be run for each sample ${\mathbf{x}}^{(i)}$ to produce (locally) equilibrated conformations. The resulting ensemble $X^{\prime}$ will be the collection of conformations from each individual trajectory. The short MD simulations in the second stage exploits the initial swarm explored by the pre-trained diffusion sampler. The sampling with neighborhood exploitation (Str2Str-NE) enjoys (1) better exploration than conventional serial MD simulations under limited computation quota, as well as (2) better exploitation than diffusion sampler alone due to the awareness of force field.
3.3 Fine-tuning of diffusion sampler
The two-stage sampling proposed above can itself serve as a good sampler. Alternatively, given the refined conformations $X^{\prime}$ as dataset, one can incorporate the supervision from the force field of MD simulation into the diffusion sampler Str2Str via fine-tuning (Str2Str-FT). In contrast to Arts et al. (2023) which involves expensive long MD trajectory for training, the acquisition of the production trajectories is very efficient: (1) The sampling (exploration) cost of diffusion sampler is negligible to MD simulations; (2) The computation of the local exploitation from short simulations can be readily paralleled according to the number of prior samples and enjoy perfect scalability.
$\mathcal{L}_{\textrm{FT}}=\mathbb{E}_{t\in[0,T]}\,\mathbb{E}_{{\mathbf{x}}_{0}\sim p({\mathbf{x}})}\left\{\omega(t,{\mathbf{x}}_{0})\,\mathbb{E}_{{\mathbf{x}}_{t}\sim p({\mathbf{x}}_{t}|{\mathbf{x}}_{0})}\left[\|{\mathbf{s}}_{\theta}({\mathbf{x}}_{t},t)-\nabla_{{\mathbf{x}}_{t}}\log p_{t|0}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})\|^{2}\right]\right\},$ (3)
where $\omega(t,{\mathbf{x}}_{0})>0$ is a positive loss reweighting function depending both on time $t$ and sample ${\mathbf{x}}_{0}$, and ${\mathbf{x}}_{t}\sim p_{t|0}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ is defined by the corresponding perturbation kernel. Among them, $\omega(t,{\mathbf{x}}_{0})$ is a natural generalization of the importance sampling (IS) scheme for time in the training of diffusion model (Song et al., 2020). The pseudo-code of the whole process is described in Algorithm 1.
Algorithm 1 Fine-tuning of a pre-trained conformation sampler.
1:Require: Target protein X, pre-trained diffusion sampler $\mathcal{S}_{\theta}$, MD simulator $\mathcal{M}$, simulation temperature $\beta$, simulation time $T^{\textrm{sim}}$.
2:$\{{\mathbf{x}}^{(i)}\}_{i=1}^{m}\sim\mathcal{S}_{\theta}(X)$ // Generate $m$ initial samples
3:$D_{X}=\varnothing$ // Initialize fine-tuning dataset
4:for $i=1$ to $m$ do
5:     $\{\hat{{\mathbf{x}}}^{(i)}_{j}\}_{j=1}^{n_{i}}\sim\mathcal{M}({\mathbf{x}}^{(i)};\beta,T^{\textrm{sim}})$ // Run short MD simulation starting from $x_{i}$
6:     $D_{X}\leftarrow D_{X}\cup\{\hat{{\mathbf{x}}}^{(i)}_{j}\}_{j=1}^{n_{i}}$ // Update dataset with the trajectory of size $n_{i}$
7:Update $\theta\rightarrow\theta^{*}$ based on $D_{X}$ by training on Eq. 3.
8:return $s_{\theta^{*}}$
4 Experiments
Setup.
To validate the effectiveness of the proposed fine-tuning pipeline, we evaluate the sampling performance on the fast-folding protein set (Lindorff-Larsen et al., 2011). For baselines, we compare the fine-tuned Str2Str (namely Str2Str-FT) against: AF2-RAVE (Vani et al., 2023), EigenFold (Jing et al., 2023), idpGAN (Janson et al., 2023), and Str2Str (pre-trained only) (Lu et al., 2024). The size of sampled ensembles is aligned to be 1,000 to make fair comparison. We follow the evaluation metrics used in Lu et al. (2024) and list them as below [^3] We omit the computationally extensive diversity metrics which involve pairwise TM-score since the distribution distance metrics (JS) can also reflect diversity mismatch.: (1) Validity: Val-Clash, Val-Bond evaluates whether the conformations obey basic geometric constraints w.r.t. steric clash and bonding association. These metrics are the higher the better ($\uparrow$). (2) Distance: JS-PwD, JS-TIC, JS-Rg assess the distance between the model distribution and reference distribution, which includes the pairwise distance (PwD), time-lagged independent components (TIC) (Naritomi & Fuchigami, 2011; Pérez-Hernández et al., 2013) and radius of gyration (Rg), which are the lower the better ($\downarrow$).
Benchmark results.
In this study, we limit the few-shot budget of MD production simulations to be 100ns in total, which is computationally tractable and lasts only several GPU hours. Under this budget, we firstly sample for each target 100 conformations from a pre-trained Str2Str followed by 100 short simulations per 1 ns long. As shown in Table 1, both the sampler with neighborhood exploitation (Str2Str-NE) and the fine-tuned one (Str2Str-FT) achieve the state-of-the-art performance across all three distance metrics while keeping good validity.
Table 1: Benchmark results of conformation sampling methods on fast folding proteins (Lindorff-Larsen et al., 2011). Metrics are averaged across all fast-folding targets for each method. The results of Str2Str is reproduced by running the open-source code, while the results of other baselines are taken from Lu et al. (2024). The best result for each section is bolded.
Methods Val-Clash($\uparrow$) Val-Bond($\uparrow$) JS-PwD($\downarrow$) JS-TIC($\downarrow$) JS-Rg ($\downarrow$)
AF2-RAVE 0.999 0.997 0.634 0.624 0.656
EigenFold 0.812 0.874 0.530 0.497 0.666
idpGAN 0.960 0.032 0.480 0.517 0.661
Str2Str 0.977 0.982 0.348 0.400 0.365
Str2Str-NE (ours) 0.990 1.000 0.294 0.369 0.331
Str2Str-FT (ours) 0.966 0.948 0.303 0.366 0.350
Ablation study.
To better demonstrate the effectiveness of the proposed Str2Str-FT, we also consider ablation experiments. The base model is the (1) Str2Str-FT sampler which is fine-tuned by samples from Str2Str-NE. We (2) replace the fine-tune data with conventional MD simulation of 100ns; and (3) (4) ablate the pre-training from both settings. The results are shown in Table 2.
Table 2: Ablation results on different strategies of training. ”NE” indicates the training set is from the samples of Str2Str-NE introduced above while ”MD” from the conventional MD simulations. ”FT” means model parameters are initialized from the PDB pre-trained checkpoint (i.e. Str2Str).
Settings PDB MD NE Val-Clash($\uparrow$) Val-Bond($\uparrow$) JS-PwD($\downarrow$) JS-TIC($\downarrow$) JS-Rg ($\downarrow$)
FT on NE data 0.966 0.948 0.303 0.366 0.350
FT on MD data 0.999 0.961 0.538 0.570 0.544
NE data 0.227 0.562 0.398 0.406 0.473
MD data 0.835 0.632 0.549 0.534 0.603
5 Conclusion
In this work, we explore the few-shot setting of protein conformation sampling by combining score-based generative models and physical MD simulations. Generative models explore the conformation space efficiently while the physical force field refine the unreasonable samples, achieving good exploration-exploitation balance. The fine-tuned diffusion sampler has shown significant improvement on conformation sampling, which exemplified the domain adaption of Str2Str with efficient short simulations. Future works can focus on the study of more effective fine-tuning strategies and enabling the sampling approximately from Boltzmann distribution.
Appendix A Experimental details
A.1 Simulation protocol
The short molecular dynamics (MD) simulations were respectively initiated from each sample from the diffusion sampler. The sampled protein models are firstly pre-processed by the PDBFixer of OpenMM (Eastman et al., 2023) before simulation. Hydrogen atoms are added under neutral environment pH=7.0. All simulations were run on the NVIDIA Tesla V100-SXM2-32GB GPU using the Amber ff14SB force field (Maier et al., 2015) and the TIP3P water model (Jorgensen et al., 1983) compatible with the corresponding Amber 14 force field. $\text{Na}^{+}$ and $\textrm{Cl}^{-}$ ions were added to achieve the electrical neutrality of the solvent. The particle mesh Ewald (PME) method (Darden et al., 1993) was used to evaluate the long-range electrostatic interaction. Each system was locally energy minimized using the L-BFGS algorithm (Liu & Nocedal, 1989) followed by equilibration in the NVT (canonical) ensemble for 1 ns. Langevin Middle Integrator (Zhang et al., 2019) with a friction coefficient of 1 $\textrm{ps}^{-1}$ was adopted as the thermostat (same below). Simulation temperatures for each benchmark target follows the original setting in Lindorff-Larsen et al. (2011), specifically, unit in Kelvin (K): Chignolin: 340K, Trp-cage: 290K, BBA: 325K, Villin: 360K, WW domain: 360K, NTL9: 355K, BBL: 298K, Protein B: 340K, Homeodomain: 360K, Protein G: 350K, $\alpha$3D: 370K and lambda-repressor: 350K. Then the system was equilibrated in the NPT (isothermal-isobaric) ensemble for 1 ns while coupled with the Monte Carlo barostat at constant 1 bar pressure. All production simulations were performed using OpenMM v8.0.0 (Eastman et al., 2023) in the NPT ensemble integrated with time step of $2.5$ fs and trajectory frames were saved per 10 ps of simulation.
A.2 Fine-tuning dataset
To acquire the fine-tuning dataset (ES) for each target, we firstly generate the seeding structure using ESMFold (Lin et al., 2023) with default configuration and ran Str2Str as described in Lu et al. (2024) to generate 100 conformations from the seeding structure. Then MD simulations was performed in parallel initialized respectively from these conformations according to the protocol above in Appendix A.1. We ignored those simulation threads failed due to possible bad initial geometries. The resulting conformations as well as their associated potential energies were saved to construct the fine-tuning dataset. We split the dataset into training and validation set with ratio $0.95:0.05$.
A.3 Setup of training and inference
To optimize the network parameters of the pre-trained Str2Str, the Adam optimizer (Kingma & Ba, 2014) was used with $\rm{lr}=10^{-4}$ and $\beta_{1}=0.9,\beta_{2}=0.999$. As scheduler, the learning rate was reduced in the factor of 0.1 when the loss has stopped decreasing for 10 epochs. The training process was set to be at least 100 epochs and mediated the early stopping strategy by monitoring the validation loss with 10 epochs patience. The fine-tuning time for different targets was on-average $\sim$ 2 GPU hours on the NVIDIA Tesla V100-SXM2-32GB GPU. For simplicity, we use uniform weighting in this study for the training objective, i.e. $\omega(x,t)\equiv 1$. Other model configurations were kept the same as in Lu et al. (2024). After fine-tuning, we sampled the target protein from the prior distribution $p_{T}({\mathbf{x}})$ with discretization of 1,000 timesteps by reversing the diffusion process via Langevin dynamics. In practice, a minimum time $t_{\epsilon}=0.01$ instead of zero is used for numerical stability. The SDE mode of sampling was used for both Str2Str and Str2Str-FT.
A.4 Fast-folding Targets.
The fast-folding targets with the reference MD trajectories are listed as follows (Lindorff-Larsen et al., 2011): Chignolin, Trp-cage, BBA, Villin, WW domain, NTL9, BBL, Protein B, Homeodomain, Protein G, $\alpha$3D and Lambda-repressor. The reference MD trajectories for evaluation were obtained by sending requests to the authors of Lindorff-Larsen et al. (2011).
A.5 Extended experimental results
Compare neighborhood exploitation with MD simulation
We demonstrate the effectiveness of the neighborhood exploitation (NE) over conventional MD simulation in Table 3. The exploration over conformation space of MD simulation (under limited budget) can be strongly enhanced by firstly sampling from a pre-trained diffusion sampler. The simulation protocol of both Str2Str-NE and MD are kept the same as described in A.1.
Table 3: Comparison between samples obtained from neighborhood exploitation (parallel short simulations) and computation-equivalent conventional MD simulation.
Setting Val-Clash($\uparrow$) Val-Bond($\uparrow$) JS-PwD($\downarrow$) JS-TIC($\downarrow$) JS-Rg ($\downarrow$)
MD (100 ns) 1.000 1.000 0.518 0.547 0.523
Str2Str-NE 0.990 1.000 0.294 0.369 0.331
Runtime Analysis
To illustrate the efficiency of proposed few-shot sampling over conventional MD simulation, we here profile the wall clock running time of each type of ”sampler” for the benchmark targets, shown in Table 3. Note that the runtime can vary from different proteins due to the different number of atoms. The profiling results are all reported as wall clock hours based on 4 NVIDIA Tesla V100-SXM2-32GB GPUs.
Figure 3: The runtime profile of different samplers used in this study

### 11 Mar 2024 v2

@@
 Molecular dynamics simulation.
-Molecular dynamics (MD) simulations amount to evolve the whole particle system through time $T^{\textrm{sim}}>0$  [^2] To avoid ambiguity with the time boundary $T$ in SGMs, the simulation time is denoted by $T^{\textrm{sim}}$ using superscript sim.
-by the physical dynamics ${d{\mathbf{x}}}={\mathbf{v}}({\mathbf{x}},\tau){d\tau}$, where the velocity field ${\mathbf{v}}\in\mathbb{R}^{N\times 3}$ can be determined by a pre-defined force field $U({\mathbf{x}})$ (aka energy or unnormalized density) via the Newtonian equation of motion: $d{\mathbf{v}}({\mathbf{x}},\tau)=\bar{{\mathbf{m}}}\nabla_{\mathbf{x}}U({\mathbf{x}})~{}d\tau$, where $\bar{{\mathbf{m}}}$ is the inverse mass of atoms in the system.
-At each time step, the position and velocity may be updated by a symplectic integrator such as Verlet (Verlet, 1967), yielding a simulation trajectory.
+Molecular dynamics (MD) simulations amount to evolve the whole particle system through time $T^{\textrm{sim}}>0$ by the physical dynamics ${d{\mathbf{x}}}={\mathbf{v}}({\mathbf{x}},\tau){d\tau}$, where ${\mathbf{v}}\in\mathbb{R}^{N\times 3}$ is the velocity field.
+With some force field defining $U({\mathbf{x}})$ as potential energy, the velocity can be updated through some dynamics such as Newtonian equation [^2] Sometimes referred to as Hamiltonian dynamics.
+Besides, stochastic dynamics can also be applied such as Langevin dynamics to update the velocity states.: $d{\mathbf{v}}({\mathbf{x}},\tau)=\bar{{\mathbf{m}}}\nabla_{\mathbf{x}}U({\mathbf{x}})~{}d\tau$, where $\bar{{\mathbf{m}}}$ is the inverse mass of atoms in the system.
+At each time step, the position and velocity can be updated by an integrator such as Verlet (Verlet, 1967), yielding a simulation trajectory.
 3.2 Neighborhood exploitation with short MD
@@
 Here arises an interesting question: how to organically combine the advantages of both generative models and physical simulations?
-Figure 2: Illustration of the conformation sampling of Str2Str (Lu et al., 2024).
-In (a), sampling is performed in two steps to obtain i.i.d.
-samples of the target protein, where the energy landscape (or information of the force field) is unknown and colored gray.
+Figure 2: Illustration of the conformation sampling scenario of Str2Str.
+In (a), sampling is performed in two steps to obtain independent samples of the target protein, where the energy landscape (or information of the force field) is unknown and colored in gray.
 In (b), a hypothetical zoom-in neighborhood of a sample is shown.
-Due to the generally complicated landscape of force field, the samples directly generated by Str2Str are probably not energy-optimal.
-Short MD simulation can be run to obtain locally equilibrium samples and can be used to fine-tune the pre-trained sampler.
+Due to the complex conformation landscape, the samples directly generated by Str2Str are probably not potential energy-optimal.
+Short MD simulation can be run to obtain locally equilibrated samples which can be used to fine-tune the pre-trained sampler.
 Exploration-exploitation balance.
@@
 To validate the effectiveness of the proposed fine-tuning pipeline, we evaluate the sampling performance on the fast-folding protein set (Lindorff-Larsen et al., 2011).
-For baselines, we compare the fine-tuned Str2Str (namely Str2Str-FT) against: AF2-RAVE (Vani et al., 2023), EigenFold (Jing et al., 2023), idpGAN (Janson et al., 2023), and Str2Str (pre-trained only) (Lu et al., 2024).
+For baselines, we compare the fine-tuned Str2Str (namely Str2Str-FT) against: MSA subsampling (Del Alamo et al., 2022), EigenFold (Jing et al., 2023), idpGAN (Janson et al., 2023), and Str2Str (pre-trained only) (Lu et al., 2024).
 The size of sampled ensembles is aligned to be 1,000 to make fair comparison.
 We follow the evaluation metrics used in Lu et al.
-(2024) and list them as below [^3] We omit the computationally extensive diversity metrics which involve pairwise TM-score since the distribution distance metrics (JS) can also reflect diversity mismatch.: (1) Validity: Val-Clash, Val-Bond evaluates whether the conformations obey basic geometric constraints w.r.t.
+(2024) and list them as below: (1) Validity: Val-Clash, Val-Bond evaluates whether the conformations obey basic geometric constraints w.r.t.
 steric clash and bonding association.
@@
 Under this budget, we firstly sample for each target 100 conformations from a pre-trained Str2Str followed by 100 short simulations per 1 ns long.
-As shown in Table 1, both the sampler with neighborhood exploitation (Str2Str-NE) and the fine-tuned one (Str2Str-FT) achieve the state-of-the-art performance across all three distance metrics while keeping good validity.
+As shown in Table 1, both the sampler with neighborhood exploitation (Str2Str-NE) and the fine-tuned version (Str2Str-FT) achieve the state-of-the-art performance across all three distance metrics while keeping good validity.
 Table 1: Benchmark results of conformation sampling methods on fast folding proteins (Lindorff-Larsen et al., 2011).
@@
 Methods Val-Clash($\uparrow$) Val-Bond($\uparrow$) JS-PwD($\downarrow$) JS-TIC($\downarrow$) JS-Rg ($\downarrow$)
-AF2-RAVE 0.999 0.997 0.634 0.624 0.656
+MSA subsampling 0.999 0.997 0.634 0.624 0.656
 EigenFold 0.812 0.874 0.530 0.497 0.666
@@
 5 Conclusion
-In this work, we explore the few-shot setting of protein conformation sampling by combining score-based generative models and physical MD simulations.
+In this work, we explore the few-shot settings of protein conformation sampling by combining score-based generative models and physical MD simulations.
 Generative models explore the conformation space efficiently while the physical force field refine the unreasonable samples, achieving good exploration-exploitation balance.
 The fine-tuned diffusion sampler has shown significant improvement on conformation sampling, which exemplified the domain adaption of Str2Str with efficient short simulations.
-Future works can focus on the study of more effective fine-tuning strategies and enabling the sampling approximately from Boltzmann distribution.
+Future works can focus on the study of more effective fine-tuning strategies to enable sampling approximately from Boltzmann distribution.
+Acknowledgments
+This project is supported by Twitter, Intel, the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund, an NRC Collaborative R&D Project (AI4D-CORE-06) as well as the IVADO Fundamental Research Project grant PRF-2019-3583139727.
 Appendix A Experimental details
@@
 The resulting conformations as well as their associated potential energies were saved to construct the fine-tuning dataset.
-We split the dataset into training and validation set with ratio $0.95:0.05$.
+We randomly split the dataset into training and validation set with ratio $0.95:0.05$ while the latter is used for early stopping of training.
 A.3 Setup of training and inference
@@
 $\omega(x,t)\equiv 1$.
-Other model configurations were kept the same as in Lu et al.
+Other model configurations and evaluation pipelines were kept the same as in Lu et al.
 (2024).
@@
 The SDE mode of sampling was used for both Str2Str and Str2Str-FT.
-A.4 Fast-folding Targets.
+A.4 Fast-folding Targets
 The fast-folding targets with the reference MD trajectories are listed as follows (Lindorff-Larsen et al., 2011): Chignolin, Trp-cage, BBA, Villin, WW domain, NTL9, BBL, Protein B, Homeodomain, Protein G, $\alpha$3D and Lambda-repressor.
@@
 (2011).
-A.5 Extended experimental results
+Appendix B Evaluation metrics
+[…]
+Appendix C Extended experimental results
 Compare neighborhood exploitation with MD simulation
@@
 Runtime Analysis
-To illustrate the efficiency of proposed few-shot sampling over conventional MD simulation, we here profile the wall clock running time of each type of ”sampler” for the benchmark targets, shown in Table 3.
+To illustrate the efficiency of proposed few-shot sampling over conventional MD simulation, we here profile the wall clock running time of each type of ”sampler” for each benchmark target, shown in Table 3.
 Note that the runtime can vary from different proteins due to the different number of atoms.
-The profiling results are all reported as wall clock hours based on 4 NVIDIA Tesla V100-SXM2-32GB GPUs.
-Figure 3: The runtime profile of different samplers used in this study
+The profiling results are all reported as the wall clock hours based on 4$\times$ NVIDIA Tesla V100-SXM2-32GB GPUs.
+Figure 3: The runtime profile of different samplers across each fast folding target used in this study.
+Appendix D Limitations and Discussion
+Similar to Str2Str (Lu et al., 2024), the proposed methods cannot guarantee the sampled ensemble are Boltzmann distributed (Noé et al., 2019) due to the nature of local simulation.
+It is not designed for replacing canonical MD simulations for the study of protein dynamics.
+Incorporating short MD simulations in the Str2Str sampler takes the advantages from both physical and neural sides: the neural sampler (eg., diffusion) is good at proposing data-like conformation hypothesis very efficiently in the vast conformation space while the physical simulation can equilibrate the samples onto some local minima.
+Both neighborhood exploitation (NE) and fine-tuning (FT) provide a simple yet effective way to refine the resulting conformations sampled from a neural sampler.
+From another perspective, neighborhood exploitation coincides with the purpose of enhanced sampling methods.
+Enhanced sampling methods aim to overcome the energy barrier and accelerate the exploration of long MD simulations.
+For example, the umbrella sampling (Torrie & Valleau, 1977), and parallel tempering (synonymous replica exchange molecular dynamics, REMD) (Hansmann, 1997; Sugita & Okamoto, 1999; Swendsen & Wang, 1986).
+The neural sampler such as Str2Str can provide diverse conformation seeds to augment the exploration of MD simulation similar to Vani et al.
+(2023).
+We find it a promising research direction to develop the neural enhanced sampling methods in the future.